OpenStudy (anonymous):

Please check. American Airline flights are on time 80% of the time. Suppose 15 flights are randomly selected, and the number of on flight times is recorded. Find the probability that at least 10 flights are on time. This is what I'm doing: P(X<10) = 15 choose 9*(0.2)^9*(0.8)^6 following the binomial probability formula, but I don't get the correct answer in the book. Please help because it's driving me nuts!

OpenStudy (anonymous):

\[P(X \le 10)\]

OpenStudy (queelius):

At least 10 flights are on time. That means it's 10 or more. So, P(X>=10) = 1 - P(X < 10) = 1 - BIN(x = 10; p=.8, n = 15).

OpenStudy (queelius):

Correction, it should be 1 - BIN(x < 10; p = .8, n = 15), where BIN is the CDF for the binomial.